OpenStudy (anonymous):

if the function p(d)=1+d/33 gives the pressure in atmospheres (atm) at the depth d in the sea (d is in feet) . note that the p(0) =1atm .p(33)=2. fwhat would be the pressure at 10 feet?

jimthompson5910 (jim_thompson5910):

\[\large p(d)=1+\frac{d}{33}\] \[\large p(10)=1+\frac{10}{33}\] \[\large p(10)=\frac{33}{33}+\frac{10}{33}\] \[\large p(10)=\frac{33+10}{33}\] \[\large p(10)=\frac{40}{33}\] \[\large p(10)\approx1.303\] So the pressure at 10 ft is approximately 1.303 atm